A tank can safely hold gas up to a pressure of 44.3 atm. When
the tank contains some nitrogen gas at 25 ºC, the gas exerts a
pressure of 10.1 atm. What is the highest temperature to
which the gas sample can be heated safely?
1
Expert's answer
2020-06-16T14:59:16-0400

According to Gay-Lussac's law, p1/T1 = p2/T2

T2 = 25oC = 298K

p1 = 44.3 atm

p2 = 10.1 atm

T1 = (p1 x T2) / p2 = (44.3atm x 298K) / 10.1atm = 1307 K

Answer: 1307 K